In how many ways can I divide a group of 12 persons (n = 12) into four subgroups (k = 4) of at least 2 persons (l=2)?
Answer by Edwin McCravy(20054)

Let the groups be group A,B,C,D

The numbers in the 4 groups can be any of 5 ways:

2,2,2,6,  that's (12C2)(10C2)(8C2)(6C6) =  66*45*28*1 =   83160
There are 4!/3! = 4 distinguishable permutations of that grouping. 
[BTW, FYI those 4 are 2,2,2,6;  2,2,6,2;   2,6,2,2; and 6,2,2,2.)  
So there are 83160*4 = 332640 for that grouping.

2,2,3,5,  that's (12C2)(10C2)(8C3)(5C5) =  66*45*56*1 =  166320 
There are 4!/2! = 12 distinguishable permutations of that grouping. 
So there are 166320*12 = 1995840 for that grouping.

2,2,4,4,  that's (12C2)(10C2)(8C4)(4C4) =  66*45*70*1 =  173250
There are 4!/(2!2!) = 6 distinguishable permutations of that grouping. 
So there are 173250*6 = 1039500 for that grouping.

2,3,3,4,  that's (12C2)(10C3)(7C3)(4C4) = 66*120*35*1 =  277200
There are 4!/(2!) = 12 distinguishable permutations of that grouping. 
So there are 277200*12 = 3326400 for that grouping.

3,3,3,3,  that's  (12C3)(9C3)(6C3)(3C3) = 220*84*20*1 =  369600
There are 4!/4! = only 1 permutation of that grouping. 
So there are 369600.

Adding those together,

332640+1995840+1039500+3326400+369600 = 7063980
---------------------------------------------------------------
Edwin